an airplane lands and starts down the runway at a northwest velocity of 55.7 m/s. what constant acceleration allows it to come to a stop in 1.24 km? (in m/s^2)
1
Expert's answer
2019-09-23T09:17:22-0400

Solution:

"\upsilon=\upsilon_0-at"

But

"\upsilon=0"

So:

"\upsilon_0=at"

And than:

"t=\frac{\upsilon_0}{a}"

Distance before the stopping:

"S=\upsilon_0t-\frac{at^2}{2}=\frac{\upsilon^2_0}{a}-\frac{a}{2}\frac{\upsilon^2_0}{a^2}=\frac{\upsilon^2_0}{2a}"


Than we can find the acceleration:

"a=\frac{\upsilon^2_0}{2S}=\frac{(55.7)^2}{2\cdot{}1.24\cdot{}10^3}=1.25" m/s2.


Answer:

a = 1.25 m/s2.